OpenStudy (anonymous):

On a biology poster, a beetle is drawn to the scale of 9 cm:2 cm. If the drawing on the poster is 28 cm long,how long is the actual beetle.

OpenStudy (campbell_st):

9/2 = 28/x then 9x = 56 or x = 56/9 arround 6.22 cm

Directrix (directrix):

9/2 = B/28 where B represents the size of the actual Beatle 2B = 9*28 B = 126 cm

OpenStudy (campbell_st):

wow a beetle 126cm long... I don't want to meet that guy